# Goal-Conditioned Data Augmentation for Offline Reinforcement Learning

### v1

@@ VII-C5 Ablation on Mixed Datasets
We further present the results for baseline data augmentation methods when combining original datasets with augmented datasets, as shown in Table VII. For Gym tasks, TATU and SynthER exhibit degraded performance when the original datasets are included during training, while their performance remains comparable on Maze2D tasks. In contrast, DStitch demonstrates the opposite trend, performing comparably on Gym tasks but worse on Maze2D tasks when using the mixed datasets. Our GODA shows reduced scores when the original datasets are removed; however, its final results remain superior to other baselines. GODA only delivers slightly worse performance than DStitch on Gym tasks when evaluated with TD3+BC but still outperforms TATU and SynthER across all tasks.
Figure 8: Ablation on mixed datasets for IQL evaluation.
TABLE VII: Data quality evaluation metrics for SynthER and GODA on Walker2D tasks. Smaller Dynamics MSE, larger L2 Distance, and larger Average Reward indicate better quality.
@@
Maze2D Average w/o Original Data 45.7±8.2 45.6±2.9 47.5±6.8 66.7±9.3 68.2±10.6 65.1±29.7 65.9±24.9 74.5±11.2
@@
VII-C6 Ablation on Size of Original Dataset
@@
Figure 9 illustrates the performance of different evaluation algorithms trained separately on the original datasets and on the mixture of original and augmented datasets. Notably, we only reduce the size of the original datasets used for directly training the evaluation algorithms or for training GODA, while maintaining the same number of augmented samples (20K) across all cases. Specifically: For the ”Original” method, the evaluation algorithms are trained solely on the reduced original datasets. For GODA, the GODA model is first trained on the reduced original datasets, and then 20K augmented samples are generated. The evaluation algorithms are subsequently trained on the mixed datasets.
@@
VIII Conclusion and Discussion
This paper proposes a novel goal-conditioned data augmentation method for offline RL, namely GODA, which integrates goal guidance into the data augmentation process. We define the easily obtainable return-to-go signal, along with its corresponding timestep in a trajectory, as the goal condition. To achieve high-return augmentation, we introduce several goal selection mechanisms and a scaling method. Additionally, we propose a novel adaptive gated conditioning structure to better incorporate goal conditions into our diffusion model. We demonstrate that data augmented by GODA shows higher quality than SynthER without goal conditions on different evaluation metrics. Extensive experiments on the D4RL benchmark confirm that GODA enhances the performance of classic offline RL methods when trained on GODA-augmented datasets. Furthermore, we evaluate GODA on real-world traffic signal control tasks. The results demonstrate that GODA is highly applicable to TSC problems, even with very small real-world training datasets, making RL-based methods more practical for real-world applications. In the future, we plan to improve the performance of GODA and explore the application of GODA on more complex real-world challeges.

### v2

@@ C.5 Ablation on Mixed Datasets
We further present the results for baseline data augmentation methods when combining original datasets with augmented datasets, as shown in Table 11. For Gym tasks, TATU and SynthER exhibit degraded performance when the original datasets are included during training, while their performance remains comparable on Maze2D tasks. In contrast, DStitch demonstrates the opposite trend, performing comparably on Gym tasks but worse on Maze2D tasks when using the mixed datasets. Our GODA shows reduced scores when the original datasets are removed; however, its final results remain superior to other baselines. GODA only delivers slightly worse performance than DStitch on Gym tasks when evaluated with TD3+BC, but still outperforms TATU and SynthER across all tasks.
Table 11: Ablation study on mixed datasets for baseline methods.
@@
Maze2D Avg. w/o Original Data 45.7±8.2 45.6±2.9 47.5±6.8 66.7±9.3 68.2±10.6 65.1±29.7 65.9±24.9 74.5±11.2
@@
C.6 Ablation on Size of Original Dataset
@@
Figure 12 illustrates the performance of different evaluation algorithms trained separately on the original datasets and on the mixture of original and augmented datasets. Notably, we only reduce the size of the original datasets used for directly training the evaluation algorithms or for training GODA, while maintaining the same number of augmented samples (20K) across all cases. Specifically, for the "Original" method, the evaluation algorithms are trained solely on the reduced original datasets. For GODA, the GODA model is first trained on the reduced original datasets, and then 20K augmented samples are generated. The evaluation algorithms are subsequently trained on the mixed datasets.